OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of one-third and passes through the point (-6, 0). A. y minus 0 equals one-third times the quantity x plus 6 B. y minus 0 equals one-third times the quantity x minus 6 C. y plus 6 equals one-third times the quantity x minus 0 D. y minus 6 equals one-third times the quantity x minus 0

OpenStudy (anonymous):

Hi! you want an equation of the form\[y = m(x - x_o)+ y_o\]

OpenStudy (anonymous):

Your know that \[m = \frac{1}{3}\]

OpenStudy (anonymous):

and \[(x_0,y_0)=(-6,0)\]

OpenStudy (anonymous):

So you just need tu plug those values into the equation, so: \[y = \frac{1}{3}(x-(-6))+0\]

OpenStudy (anonymous):

Simplifying you get: \[y = \frac{1}{3}x+2\]

OpenStudy (anonymous):

none of those answers are on here

OpenStudy (anonymous):

i'm so confused

OpenStudy (anonymous):

wait a minute.

OpenStudy (anonymous):

Look carefully, the answer is there.

OpenStudy (anonymous):

it doesn't have that available in the answer choices

OpenStudy (anonymous):

yes it is available.

OpenStudy (anonymous):

okay then , i'll type these in the equation form , you tell me which answer it is .

OpenStudy (anonymous):

You can write the last equation I wrote like this: \[y = \frac{1}{3}(x+6)\]

OpenStudy (anonymous):

Is that equation available at the options?

OpenStudy (anonymous):

yeah , you forgot the y - 0 = though . okay now i see it . thanks .

OpenStudy (anonymous):

well is not necessary to write the 0

OpenStudy (anonymous):

You're welcome.